# VectorFit : Adaptive Singular & Bias Vector Fine-Tuning of Pre-trained Foundation Models

### v1

@@ Abstract
Popular PEFT methods achieve parameter efficiency by assuming that incremental weight updates are inherently low-rank, which often leads to a performance gap compared to full fine-tuning. While recent methods have attempted to address this limitation, they typically lack sufficient parameter and memory efficiency. We propose VectorFit, an effective and easily deployable approach that adaptively trains the singular vectors and biases of pre-trained weight matrices. We demonstrate that the utilization of structural and transformational characteristics of pre-trained weights enables high-rank updates comparable to those of full fine-tuning. As a result, VectorFit achieves superior performance with $\boldsymbol{\times}$ less trainable parameters compared to state-of-the-art PEFT methods. Through extensive experiments over 17 datasets spanning diverse language and vision tasks such as natural language understanding and generation, question answering, image classification, and image generation, we exhibit that VectorFit consistently outperforms baselines, even in extremely low-budget scenarios.
Machine Learning, ICML
@@ 1 Introduction
Pre-trained foundation models (PFMs) have set unprecedented standards in language, vision, and audio tasks (Touvron et al., 2023; Rombach et al., 2021; Radford et al., 2022), showcasing their strong performance across diverse domains. Refining these models through fine-tuning is a powerful approach to enhance their performance across diverse downstream tasks (Li et al., 2023; Yu et al., 2024). This process helps models adhere to given instructions (Xu et al., 2024a), adopt preferred behaviors, and discard undesirable ones (Rafailov et al., 2023). However, adapting these models to downstream tasks through full fine-tuning (Full-FT) is a significant challenge, primarily due to the immense computational and memory overhead. For instance, models like DeBERTa-V3 (He et al., 2023) with 300 million parameters, ViT-22B (Dehghani et al., 2023) with 22 billion parameters, and Llama-3 (Grattafiori et al., 2024) with a staggering 405 billion parameters exemplify the scale of modern PFMs. Adapting these models for multiple downstream tasks is resource-intensive and typically requires maintaining separate copies of the full model for each task, leading to a very high memory consumption.
@@
Parameter-Efficient Fine-Tuning (PEFT) mitigates these challenges by introducing a small set of trainable parameters to produce specialized models. For example, PFMs like LLMs are fine-tuned for tasks such as text classification (Wang et al., 2019), question answering (Rajpurkar et al., 2016), and text generation (Li et al., 2021) using task-specific datasets. PEFT techniques, such as LoRA (Hu et al., 2021) and adapter (Pfeiffer et al., 2021), significantly reduce the number of trainable parameters compared to Full-FT, although this can sometimes compromise performance. More advanced methods like AdaLoRA (Zhang et al., 2023b) try to increase expressiveness by adaptively choosing trainable parameters, bridging the performance gap. However, majority of successful PEFT methods rely on the assumption that incremental weight matrices are low-rank, which limits their expressiveness and disregards the nuances of weight matrix transformation during Full-FT. Moreover, even state-of-the-art PEFT techniques (e.g., LoRA, Adapter, and AdaLoRA) can still result in a substantial number of trainable parameters, even in their highest parameter-efficient setup (e.g., LoRA with rank 1). Although there are a few recent methods that do not rely on low-rank updates (Qiu et al., 2023; Lingam et al., 2024), all of them, to the best of our knowledge, work based on fine-tuning a newly initialized set of weight matrices. This leads to a high overall parameter count and memory consumption, twice as that of LoRA in SVFT (Lingam et al., 2024).
@@
This prompts the question: Can we achieve high-rank adaptation without incurring prohibitive parameter and memory costs? As an answer, we introduce VectorFit, which directly leverages the structural and transformational characteristics of the pre-trained weights instead of introducing new weights for fine-tuning, differentiating our method from prior work. Given a pre-trained weight matrix $W_{0}$, VectorFit applies singular value decomposition (SVD), such that $W_{0}=U\Sigma V^{T}$. The method then selectively adapts the singular vector ($\Sigma$) and the bias ($b$) associated with $W_{0}$, focusing on those $\Sigma$ and $b$ that exhibit suboptimal training compared to those of other weight matrices. We propose a mechanism called Adaptive Vector Freezing to achieve this.
Since the singular vectors represent the stretching of the weight matrices in their high-dimensional subspace, directly fine-tuning them allows for high expressiveness (Appendix D.5). Additionally, training the bias vectors gives translational degree of freedom, further enhancing the expressiveness during fine-tuning.
@@
Hence, VectorFit performs high-rank updates comparable to Full-FT (Figure 9) while using significantly fewer trainable parameters ($\leq 0.1\%$). It outperforms the baselines in terms of performance relative to parameter efficiency (Figure 1). VectorFit also shows a practical memory consumption approximately equivalent to that of LoRA with rank 1 (Figure 5).
@@ 2 Related Work
Adapter-Based methods. This research direction emphasizes incorporating small neural networks into PFMs and fine-tuning only these modules for specific tasks, keeping the base model frozen and shared across tasks. This approach introduces a limited number of task-specific parameters, significantly improving the scalability and practicality of large models. For instance, adapter tuning (Houlsby et al., 2019; Pfeiffer et al., 2021; He et al., 2022) integrates small neural networks, known as adapters, between the layers of the base model. Other methods, such as prefix tuning (Li & Liang, 2021) and prompt tuning (Lester et al., 2021), add trainable prefix tokens to the input or hidden layers of the model. These techniques claim to have demonstrated performance comparable to Full-FT while updating less than 1% of the model parameters, significantly reducing memory requirements.
@@
LoRA-Based methods. A significant advancement in PEFT is Low-Rank Adaptation (LoRA) (Hu et al., 2021), which preserves the pre-trained model weights and incorporates trainable low-rank matrices within each transformer layer. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d_{r}\times d_{c}}$, LoRA constrains the weight update $\Delta W$ to a low-rank decomposition: $y=W_{0}x+\Delta Wx=W_{0}x+BAx$, where $B\in\mathbb{R}^{d_{r}\times r}$, $A\in\mathbb{R}^{r\times d_{c}}$ and rank $r<<min(d_{r},d_{c})$. Only $A$ and $B$ are trainable parameters.
@@
Several studies have introduced variations of the LoRA algorithm, focusing on reducing the number of trainable parameters (Zhang et al., 2024a; Kopiczko et al., 2024; Ding et al., 2023), improving the flexibility of low-rank structures (Koohpayegani et al., 2024; Zi et al., 2024; Sun et al., 2024), enabling adaptive parameter allocation (Zhang et al., 2023a), and integrating LoRA with techniques like quantization (Dettmers et al., 2024; Xu et al., 2024b) and pruning (Zhang et al., 2024b).
@@
A significant enhancement over LoRA is AdaLoRA (Zhang et al., 2023b), which addresses LoRA’s limitation of evenly distributing trainable parameters across weight matrices, ignoring their varying importance. In AdaLoRA, the incremental updates are parameterized as singular value decomposed matrices $P\Lambda Q$, where $P\in\mathbb{R}^{d_{r}\times r}$ and $Q\in\mathbb{R}^{r\times d_{c}}$ are the left and right singular matrices, $\Lambda\in\mathbb{R}^{r\times 1}$ is the singular vector. The orthogonality of $P$ and $Q$ is maintained using the regularizer $R(P,Q)=\|P^{\top}P-I\|^{2}_{F}+\|QQ^{\top}-I\|^{2}_{F}$. The rank of low-rank updates is dynamically adjusted using an importance metric derived from $\Lambda$. By pruning less significant singular values while allowing for recovery, AdaLoRA claims to have improved performance with a similar parameter budget as LoRA.
@@
Other PEFT methods. Orthogonal Fine-Tuning (OFT) (Qiu et al., 2023) introduces an orthogonal projection approach using orthogonal regularization. It focuses on optimizing parameters while preserving the orthogonality of weight updates, ensuring minimal interference with pre-trained knowledge. However, it still demands a significant number of trainable parameters because of the high dimensionality of the matrices. Butterfly Orthogonal Fine-Tuning (BOFT) (Liu et al., 2024) builds upon OFT by introducing Butterfly factorization and claims to improve parameter efficiency, and fine-tuning flexibility. Singular Vectors guided Fine-Tuning (SVFT) (Lingam et al., 2024) leverages the singular value decomposition of pre-trained weight matrices to parameterize weight updates as $y=W_{0}x+\Delta Wx=U(\Sigma+M)V^{T}x$, where $M$ is a sparse trainable matrix with pre-determined and fixed sparsity pattern. As $M$ is not restricted to be low-rank, SVFT claims to achieve high-rank gradient updates. Nonetheless, SVFT uses four matrices, $U,\Sigma,V$, and $M$, for every pre-trained weight matrix. Also, their dimensions are comparable to those of the pre-trained weight matrix. This leads to a high parameter and memory cost. On the other hand, VectorFit is an SVD-based PEFT method that directly trains the singular and bias vectors in a selective manner, significantly reducing the overall parameter and memory cost.
@@ 3.1 Vector Fine-Tuning
$y=\left(U{\color[rgb]{1.00,0.00,0.00}\definecolor[named]{pgfstrokecolor}{rgb}{1.00,0.00,0.00}\Sigma}V^{T}\right)x+{\color[rgb]{1.00,0.00,0.00}\definecolor[named]{pgfstrokecolor}{rgb}{1.00,0.00,0.00}b_{0}}$ (1)
@@ 3.2 Adaptive Vector Freezing
In Section 6.1, we theoretically and experimentally show that this mechanism leads to similar results as that of dropout. It is important to note that using the standard dropout algorithm for singular vectors results in a significant performance drop, even with a very low dropout probability. This signifies that certain singular values, and their corresponding left and right singular directions are extremely important that they cannot be dropped. Therefore, the AVF mechanism is crucial for maintaining effective training.
@@ 4 Experiments
Implementation details. All algorithms are implemented using PyTorch (Paszke et al., 2019). Our implementation builds upon the publicly available Huggingface Transformers codebase (Wolf et al., 2020). Appendix C contains the full details of our experimental setup and hyperparameter configurations.
@@ 4.1 Tasks and Datasets
    Natural Language Understanding (NLU): Experiments are conducted on the GLUE benchmark (Wang et al., 2019), which includes single-sentence classification, similarity/paraphrase, and natural language inference tasks.
@@
    Question Answering (QA): Performance is tested on SQuAD v1.1 and SQuAD v2.0 (Rajpurkar et al., 2016), treating QA as a sequence labeling problem to predict the start and end token probabilities for answer spans.
@@
    Natural Language Generation (NLG): Evaluation is performed on the XSum (Narayan et al., 2018) and CNN/DailyMail (Nallapati et al., 2016) datasets for text summarization task.
@@
    Image Classification: Our method is assessed on image classification tasks using CIFAR10 (Krizhevsky, 2009), GTSRB (Stallkamp et al., 2012), MNIST (Deng, 2012), and RESISC45 (Cheng et al., 2017) datasets.
@@
    Image Generation: The results on subject-driven image generation is evaluated using the Dreambooth dataset (Ruiz et al., 2023).
@@ 4.2 Base Models
We use four distinct base model types that are representative of a wide range of PFMs for the evaluation of our algorithm:
@@
    DeBERTaV3-base (He et al., 2023), a transformer encoder-only language model, applied to NLU and QA tasks.
@@
    BART-large (Lewis et al., 2020), a transformer encoder-decoder model, used for NLG tasks.
@@
    ViT-base (Dosovitskiy et al., 2021), a vision transformer model, applied to image classification tasks, pre-trained on Imagenet-1K.
@@
    Stable Diffusion v1.4 (Rombach et al., 2021), a latent diffusion model based on UNet architecture, used for text-to-image generation.
Baselines. We compare our approach against Full-FT, which updates all parameters across all layers. Additionally, we evaluate it against state-of-the-art methods from each of the three categories mentioned in Section 2. These include LoRA (Hu et al., 2021), AdaLoRA (Zhang et al., 2023b), PAdaptor (Pfeiffer et al., 2021), HAdaptor (Houlsby et al., 2019), and SVFT (Lingam et al., 2024).
@@ 5.2 Question Answering
Table 4: The performance results for ViT-base fine-tuned on the CIFAR10, GTSRB, MNIST, and RESISC45 datasets are presented. The # Params column indicates the proportion of trainable parameters, and the corresponding image classification accuracies are reported.
@@ 5.3 Natural Language Generation
We evaluate VectorFit on the XSum and CNN/DailyMail datasets using the ROUGE (1/2/L) metrics in Table 3. Despite a 33.3% higher relative parameter efficiency than the baselines, VectorFit consistently outperforms them on both datasets. Notably, VectorFit recovers 95% of Full-FT Rouge-L score with only 0.12% trainable parameters, compared to the baselines that recover 86% accuracy with 0.16% trainable parameters on the Xsum dataset. VectorFit’s performance surpasses Full-FT on CNN/Dailymail dataset, demonstrating superior efficiency and performance on complex tasks.

### v2

@@ Abstract
Popular PEFT methods reduce trainable parameter count for fine-tuning by parameterizing new low-rank or sparse trainable weights in parallel to the frozen pre-trained weights $W$. However, these weights are trained from scratch, and there exists a performance gap between these methods and full fine-tuning, especially in low-budget settings. We introduce VectorFit, a new way of parameterization that efficiently utilizes the existing knowledge embedded in $W$ by adaptively training their singular vectors and biases. We show that utilizing the structural and transformational properties of $W$ in this way can lead to high-rank incremental weight matrices $\Delta W$, comparable to that of full fine-tuning. VectorFit delivers superior results with $\boldsymbol{\times}$ fewer trainable parameters than the leading PEFT methods. Through comprehensive experiments across 19 datasets covering a wide range of language and vision tasks such as natural language understanding and generation, question answering, image classification, and image generation, we demonstrate that VectorFit surpasses baselines in terms of performance as a function of parameter-efficiency.
[…]
123
@@ 1 Introduction
Pre-trained foundation models (PFMs) have set unprecedented standards in language, vision, and audio tasks [39, 33, 30], showcasing their strong performance across diverse domains. Refining these models through fine-tuning is a powerful approach to enhance their performance across diverse downstream tasks [20, 45]. This process helps models adhere to given instructions [43], adopt preferred behaviors, and discard undesirable ones [31]. However, adapting these models to downstream tasks through full fine-tuning (Full-FT) is a significant challenge, primarily due to the immense computational and memory overhead. For instance, models like DeBERTa-V3 [10] with 300 million parameters, ViT-22B [3] with 22 billion parameters, and Llama-3 [8] with a staggering 405 billion parameters exemplify the scale of modern PFMs. Adapting these models for multiple downstream tasks is resource-intensive and typically requires maintaining separate copies of the full model for each task, leading to a very high memory consumption.
@@
Parameter-Efficient Fine-Tuning (PEFT) mitigates these challenges by introducing a small set of trainable parameters to produce specialized models. For example, PFMs like LLMs are fine-tuned for tasks such as text classification [40], question answering [32], and text generation [19] using task-specific datasets. PEFT techniques, such as LoRA [13] and adapter [28], significantly reduce the number of trainable parameters compared to Full-FT, although this can compromise the performance. More advanced methods like AdaLoRA [49] try to increase expressiveness by adaptively choosing trainable parameters, bridging the performance gap. However, majority of successful PEFT methods work by adding a new set of weight matrices with the assumption that incremental weight matrices are low-rank. This may limit their expressiveness and disregards the nuances of weight matrix transformation during Full-FT. Moreover, even state-of-the-art PEFT techniques (e.g., LoRA, Adapter, and AdaLoRA) can still result in a substantial number of trainable parameters, even in their highest parameter-efficient setup (e.g., LoRA with rank 1). Although there are a few recent methods that do not rely on low-rank updates [29, 22], all of them, to the best of our knowledge, work based on fine-tuning a newly initialized set of weight matrices. This leads to a high overall parameter count and memory consumption, twice as that of LoRA in SVFT [22].
@@
This prompts the question: Can we achieve extreme parameter-efficiency through the efficient use of pre-training knowledge embedded in the weight matrices, without incurring prohibitive parameter and memory costs? As an answer, we introduce VectorFit, which directly leverages the structural and transformational characteristics of the pre-trained weights instead of introducing new weights for fine-tuning, differentiating our method from prior work. Given a pre-trained weight matrix $W_{0}$, VectorFit applies singular value decomposition (SVD), such that $W_{0}=U\Sigma V^{T}$. The method then selectively adapts the singular vector ($\Sigma$) and the bias ($b$) associated with $W_{0}$, focusing on those $\Sigma$ and $b$ that exhibit suboptimal training compared to those of other weight matrices. We propose a mechanism called Adaptive Vector Freezing to achieve this.
@@
Since the singular vectors represent the stretching of the transformation applied by the weight matrices in their high-dimensional subspace, directly fine-tuning them allows for high expressiveness (Appendix D.5). This is evident from the fact that VectorFit performs high-rank updates comparable to Full-FT (Figure 10) while using significantly fewer trainable parameters ($\leq 0.1\%$). Additionally, training the bias vectors gives translational degree of freedom, further enhancing the expressiveness during fine-tuning. VectorFit outperforms the baselines in terms of performance relative to parameter efficiency (Figure 1). It also gives a practical memory consumption approximately equivalent to that of LoRA with rank 1 for smaller base models (Figure 5). The memory consumption of VectorFit lies in between LoRA and SVFT for larger base models (Figure 6).
@@ 2 Related Work
Adapter-Based methods. This research direction emphasizes incorporating small neural networks into PFMs and fine-tuning only these modules for specific tasks, keeping the base model frozen and shared across tasks. This approach introduces a limited number of task-specific parameters, significantly improving the scalability and practicality of large models. For instance, adapter tuning [12, 28, 9] integrates small neural networks, known as adapters, between the layers of the base model. Other methods, such as prefix tuning [21] and prompt tuning [17], add trainable prefix tokens to the input or hidden layers of the model. These techniques claim to have demonstrated performance comparable to Full-FT while updating less than 1% of the model parameters, significantly reducing memory requirements.
@@
LoRA-Based methods. A significant advancement in PEFT is Low-Rank Adaptation (LoRA) [13], which preserves the pre-trained model weights and incorporates trainable low-rank matrices within each transformer layer. For a pre-trained weight matrix $W_{0}\in\mathbb{R}^{d_{r}\times d_{c}}$, LoRA constrains the weight update $\Delta W$ to a low-rank decomposition: $y=W_{0}x+\Delta Wx=W_{0}x+BAx$, where $B\in\mathbb{R}^{d_{r}\times r}$, $A\in\mathbb{R}^{r\times d_{c}}$ and rank $r<<min(d_{r},d_{c})$. Only $A$ and $B$ are trainable parameters.
@@
Several studies have introduced variations of the LoRA algorithm, focusing on reducing the number of trainable parameters [47, 15, 6], improving the flexibility of low-rank structures [14, 50, 37], enabling adaptive parameter allocation [46], and integrating LoRA with techniques like quantization [5, 44] and pruning [48].
@@
A significant enhancement over LoRA is AdaLoRA [49], which addresses LoRA’s limitation of evenly distributing trainable parameters across weight matrices, ignoring their varying importance. In AdaLoRA, the incremental updates are parameterized as singular value decomposed matrices $P\Lambda Q$, where $P\in\mathbb{R}^{d_{r}\times r}$ and $Q\in\mathbb{R}^{r\times d_{c}}$ are the left and right singular matrices, $\Lambda\in\mathbb{R}^{r\times 1}$ is the singular vector. The orthogonality of $P$ and $Q$ is maintained using the regularizer $R(P,Q)=\|P^{\top}P-I\|^{2}_{F}+\|QQ^{\top}-I\|^{2}_{F}$. The rank of low-rank updates is dynamically adjusted using an importance metric derived from $\Lambda$. By pruning less significant singular values while allowing for recovery, AdaLoRA claims to have improved performance with a similar parameter budget as LoRA. Pissa [24] is another method that works similar to LoRA while the initialization of the low-rank trainable weights is done using the principal singular values and vectors.
@@
Other PEFT methods. Orthogonal Fine-Tuning (OFT) [29] introduces an orthogonal projection approach using orthogonal regularization. It focuses on optimizing parameters while preserving the orthogonality of weight updates, ensuring minimal interference with pre-trained knowledge. However, it still demands a significant number of trainable parameters because of the high dimensionality of the matrices. Butterfly Orthogonal Fine-Tuning (BOFT) [23] builds upon OFT by introducing Butterfly factorization and claims to improve parameter efficiency, and fine-tuning flexibility. Singular Vectors guided Fine-Tuning (SVFT) [22] leverages the singular value decomposition of pre-trained weight matrices to parameterize weight updates as $y=W_{0}x+\Delta Wx=U(\Sigma+M)V^{T}x$, where $M$ is a sparse trainable matrix with pre-determined and fixed sparsity pattern. As $M$ is not restricted to be low-rank, SVFT claims to achieve high-rank gradient updates. Nonetheless, SVFT uses four matrices, $U,\Sigma,V$, and $M$, for every pre-trained weight matrix. Also, their dimensions are comparable to those of the pre-trained weight matrix. This leads to a high parameter and memory cost.
@@ 3.1 Vector Fine-Tuning
$y=\left(U{\color[rgb]{1.00,0.00,0.00}\Sigma}V^{T}\right)x+{\color[rgb]{1.00,0.00,0.00}b_{0}}$ (1)
@@ 3.2 Adaptive Vector Freezing
In Section 6.1, we theoretically and experimentally show that this mechanism leads to similar effect as that of dropout. It is important to note that using the standard dropout algorithm for singular vectors results in a significant performance drop, even with a very low dropout probability. This signifies that certain singular values, and their corresponding left and right singular directions are extremely important that they cannot be dropped. Therefore, the AVF mechanism is crucial for maintaining effective training.
@@ 4 Experiments
Implementation details. All algorithms are implemented using PyTorch [27]. Our implementation builds upon the publicly available Huggingface Transformers codebase [42]. Appendix C contains the full details of our experimental setup and hyperparameter configurations.
@@ 4.1 Tasks and Datasets
    Natural Language Understanding (NLU): Experiments are conducted on the GLUE benchmark [40], which includes single-sentence classification, similarity/paraphrase, and natural language inference tasks.
@@
    Question Answering (QA): Performance is tested on SQuAD v1.1 and SQuAD v2.0 [32], treating QA as a sequence labeling problem to predict the start and end token probabilities for answer spans.
@@
    Natural Language Generation (NLG): Evaluation is performed on the XSum [26] and CNN/DailyMail [25] datasets for text summarization task. GSM8K [2] and Math [11] datasets are used to evaluate mathematical problem-solving capabilities.
@@
    Image Classification: Our method is assessed on image classification tasks using CIFAR10 [16], GTSRB [36], MNIST [4], and RESISC45 [1] datasets.
@@
    Image Generation: The results on subject-driven image generation is evaluated using the Dreambooth dataset [34].
@@ 4.2 Base Models
We use six distinct base model types that are representative of a wide range of PFMs for the evaluation of our algorithm:
@@
    DeBERTaV3-base [10], a transformer encoder-only language model, applied to NLU and QA tasks.
@@
    BART-large [18], a transformer encoder-decoder model, used for NLG tasks.
@@
    Gemma-7B [38], a transformer decoder-only model, used for NLG tasks.
@@
    Llama-3-8B [8], a transformer decoder-only model, used for NLG tasks.
  5. 5.
    ViT-base [7], a vision transformer model, applied to image classification tasks, pre-trained on Imagenet-1K.
  6. 6.
    Stable Diffusion v1.4 [33], a latent diffusion model based on UNet architecture, used for text-to-image generation.
Baselines. We compare our approach against Full-FT, which updates all parameters across all layers. Additionally, we evaluate it against state-of-the-art methods from each of the three categories mentioned in Section 2. These include LoRA [13], AdaLoRA [49], PAdaptor [28], HAdaptor [12], Pissa [24], and SVFT [22].
@@ 5.2 Question Answering
Table 4: Comparison of various methods for Gemma-7B and Llama-3-8B models, fine-tuned for mathematical reasoning. The table includes number of parameters and accuracy on GSM-8k and MATH benchmarks.
Method Gemma-7B Llama-3-8B
# Params GSM-8k MATH # Params GSM-8k MATH
Full FT 8.5B 74.67 25.70 8.0B 64.13 16.24
LoRA (r=1) 0.82M 72.4 26.28 1.77M 68.84 20.94
AdaLoRA (r=1) 0.81M 72.5 26.41 1.77M 68.73 21.09
Pissa 0.82M 72.3 26.31 1.77M 67.92 21.01
SVFT 0.43M 73.50 27.30 0.48M 69.22 20.44
VectorFit 0.43M 73.94 27.41 0.48M 70.83 21.02
Table 5: The performance results for ViT-base fine-tuned on the CIFAR10, GTSRB, MNIST, and RESISC45 datasets are presented. The # Params column indicates the proportion of trainable parameters, and the corresponding image classification accuracies are reported. Section 6.3 contains more details about the variations of VectorFit.
@@ 5.3 Natural Language Generation
We evaluate VectorFit on the XSum and CNN/DailyMail datasets using the ROUGE (1/2/L) metrics in Table 3. Despite a 33.3% higher relative parameter efficiency than the baselines, VectorFit consistently outperforms them on both datasets. Notably, VectorFit recovers 95% of Full-FT Rouge-L score with only 0.12% trainable parameters, compared to the baselines that recover 86% accuracy with 0.16% trainable parameters on the Xsum dataset. VectorFit’s performance surpasses Full-FT on CNN/Dailymail dataset, demonstrating superior efficiency and performance on complex tasks. Additionally, we evaluate VectorFit for mathematical reasoning with Gemma-7B and Llama-3-8B, as shown in Table 4, highlighting that our method scales well to larger base model sizes.